The common Eastern mole, a mammal, typically has a mass of 78 g, which corresponds to about 7.8 moles of atoms. (A mole of atoms is 6.02 × 10^23 atoms. In atomic mass units (u), what is the average mass of atoms in the common Eastern mole?
1
Expert's answer
2015-03-26T10:52:25-0400
1 atomic mass unit is 1.66053892*10^-27kg.
The 1 mole of atom weights 1.66053892*10^-27kg*6.02 × 10^23 atoms = 9.9964*10^-4kg.
The average mass of atoms in the common Eastern mole: 9.9964*10^-4kg*7.8moles= 77.97192*10^-4kg.